### v1

@@ 5.1 Baselines
To our best knowledge, the most related works to ours are NSM [67] and SAMP [25] since they both synthesize human-object interactions. We train each baseline method on the FORCE dataset using their publicly released codes. These methods neglect the physics attributes of the interaction. To measure the effectiveness of our intuitive physics encoding, we adapt these baselines by additionally conditioning on the resistive force and the binary hand contact. We refer to these modified baselines as NSM+ and SAMP+. Moreover, we evaluate two ablative variations of FORCE . One model is without the physics encoding $\mathcal{F}$, and the other one is without the contact prediction.
@@ 5.2 Quantitative Evaluation
We quantitatively evaluate our model performance with two setups. a) We evaluate the accuracy (MPJPE) of motion synthesis on testing sequences following our training/testing split explained in Section 3. The model synthesizes poses auto-regressively given the trajectories. The sequences are each distinctive, covering diverse pushing, pulling and carrying motions at different resistances. b) We evaluate all other metrics on the complete synthesis of approaching and interacting in a random setup. For fairness, the object shapes are sampled. They are placed randomly relative to the human (within 6 meters), and given a random y-rotation. The synthesis is performed 120 times, repeating each action type (pushing, pulling, one-handed, and two-handed carrying) 30 times.
MPJPE(cm) [55, 35]. The average per-joint error in centimeters, measuring the accuracy of human motion synthesis. Note, the result on “Ours no Contact” is omitted because the ablative baseline utilizes the same MNet as “Ours”.
@@
Execution time(s) [27, 53]. Average time in seconds till a successful interaction.
Diversity [24, 87, 93, 25]. We compute the Average Pairwise Distance (APD) on the root-relative joint positions of the synthesized motion. Diversity is compared against the training data.
@@
Foot-slide(cm) [31, 67, 28, 25, 91, 92]. The average distance traveled by the pivotal foot when in contact with the ground at each footstep.
Table 2: Quantitative comparisons with baselines. MPJPE is measured on the testing sequences, and the other metrics are evaluated on 120 syntheses performed with randomly sampled objects. Note, diversity on the dataset is 0.906.
@@
NSM [67] 7.71 75.1 5.94 0.707 64.3 4.59
SAMP [25] 8.67 66.2 7.72 0.719 57.7 5.79
NSM+ 7.08 79.6 6.01 0.785 69.6 4.71
SAMP+ 8.34 68.7 7.56 0.776 58.3 6.02
Ours no Phys. Enc. $\mathcal{F}$ 6.75 86.7 5.81 0.847 75.4 4.60
Ours no Contact Pred. - 94.2 5.50 0.886 77.2 4.48
Ours 6.02 97.5 5.42 0.891 84.0 4.46
Figure 4: Comparison with baselines. Our intuitive physics encoding $\mathcal{F}$ disambiguates different multi-class motions and enables the synthesis of the desired interaction (carrying in a), pulling in b). c) and d) show the coupled encoding of human force and resistance allows the synthesis of visually plausible interactions. In c), when the resistance is greater than the human force, FORCE synthesizes a more plausible “infeasible” interaction. In d), with low resistance, and the human pose maintains an upright position.
From Table 2, it can be seen that  FORCE attains the lowest MPJPE of $6.02$ cm, surpassing both baseline models and ablative variations of FORCE . This emphasizes the crucial role of our intuitive physics encoding in facilitating the learning of multi-class motions.
Under the simulated object setup, FORCE has the highest success rate of interaction of $97.5\%$ and the shortest average execution time of $5.42$ seconds, illustrating the better controllability of our approach.
Our model synthesizes motion of diversity (0.891) which is the closest match to the training data distribution (0.906). This underscores the significance of the novel physics encoding in enabling the synthesis of diverse, nuanced interactions. Additionally, our method minimizes object interpenetration, as reflected in the highest collision scores. It is notable that, compared with the ablative baseline without contact prediction, our method has a higher collision score. This indicates that with contact prediction, our model synthesizes more physically plausible motion. Furthermore, our model demonstrates improved motion realism, evident in its reduced foot-slide.
@@ 5.3 Qualitative Evaluation
Figure 5: Nuanced interactions with the same object shape, but different resistance. a) The human struggles with the interaction as the resistance increases from 50 N to 200 N. At 500 N, the human fails to interact with the object. b) Plot depicts the horizontal shift of the center of mass of the human, estimated following [73, 12]. Results averaged over 1000 frames.
@@ 5.4 Ablations
Figure 7: Without geometry (left) vs. FORCE (right). With geometry representation $\mathbf{G}$, reasons about collision and to avoid interpenetration.
@@
Figure 8: a) Different hand contacts predicted by FORCE. Contact is dependent on the resistance. b) Without contact prediction (left) vs. FORCE (right). FORCE synthesizes more physically plausible motion.
5.5 Generalization to different shapes and locations.
As illustrated in Figure 9, FORCE generalizes to different shapes. Thanks to the object shape augmentation introduced in Section 3, our model is able to synthesize lifting a chair, an object category that is not seen during training of carrying sequences.
Figure 9: (Top) FORCE generalizes to unseen shapes, demonstrated via pushing, carrying, and pulling. (Bottom) FORCE generalizes to different locations. one-handed and two-handed carry from the table and from the floor.
It is also highlighted in Figure 9 that FORCE interacts with objects at different locations, for example, picking up from different heights.
@@ 6 Conclusion
This work, FORCE , tackles the problem of synthesizing human object interactions with nuanced details, by modeling intuitive physics. It solves two main challenges. First, it proposes the first kinematic-based method that synthesizes human-object interaction conditioned on physical attributes such as object resistance and the human force. By leveraging the novel physics encoding, our method generates a diverse spectrum of interaction motions under varying resistance. Its performance in diversity and realism surpasses previous methods. Second, the accompanying FORCE dataset, featuring over 450 motion sequences, provides a valuable resource for training and evaluating kinematic-based methods.
@@ Acknowledgement
1 List of Symbols
Acknowledging the abundance of symbols in this paper, we aim to clarify and eliminate any confusion by presenting a detailed list of symbols (referenced in Table 1 List of Symbols) used in this paper, along with their corresponding definitions.
Notation
Definition
$\mathcal{F}$
intuitive physics encoding
$\mathbf{R}$
magnitude of the resistance
$\mathbf{F}$
3D human force
$\mathbf{c}$
binary desired hand contact
$\mathcal{H}$
human motion
$\mathbf{J}$
human pose
$\boldsymbol{j}^{p}_{i}$
joint position (root-relative)
$\boldsymbol{j}^{v}_{i}$
joint velocity (root-relative)
$\boldsymbol{j}^{r}_{i}$
joint rotation (root-relative)
$\mathbf{T}$
human trajectory
$\boldsymbol{t}^{p}_{i}$
ground-projected root position
$\boldsymbol{t}^{d}_{i}$
ground-projected root rotation
$\boldsymbol{t}^{a}_{i}$
current action
$\mathbf{O}$
object
$\boldsymbol{o}^{p}_{i}$
object position (root-relative)
$\boldsymbol{o}^{d}_{i}$
object rotation (root-relative)
$\boldsymbol{o}^{a}_{i}$
binary desired action
$\mathbf{G}$
geometry representation computed from the object voxels
$\phi$
phase encoding of the human motion
$\mathcal{C}$
hand contact positions
Table 3: Legend of Notations.
@@ 2 Dataset
Further details on the dataset: Our dataset comprises 450 motion sequences involving human-object interactions with a diverse range of resistance forces. Table 4 provides a detailed breakdown of the dataset categorized by the level of resistance. In this context, resistance is measured solely by the mass of the removable weight used. The masses of the objects themselves are not factored into these measurements. They are measured and will be provided with the dataset. The dataset distributions based on the type of action (Table 6), and the type of hand contact (Table 6) are also presented.
@@
Table 6: Distribution of the dataset with different hand contact.
@@
Table 6: Distribution of the dataset with different hand contact.
@@ 3 Architecture and Training Details
4 Ethics, Bias, and Boarder Impact
Our dataset was captured with the full consent and cooperation of the participant, ensuring ethical considerations were prioritized throughout the process. To mitigate potential biases, deliberate steps were taken. First, we captured objects of a range of shapes and sizes, and performed object shape augmentation to avoid biases arising from over- or under-representation of certain object shapes. Secondly, we carefully considered the distribution of resistance levels to prevent bias towards specific interaction intensities, ensuring a balanced representation of challenging and less challenging scenarios. Environmental factors, such as floor surface and lighting conditions, were also meticulously controlled and maintained constant throughout the data capture process to minimize any confounding variables.
Our dataset addresses the scarcity of accurate 3D human-object interaction data. It is the first dataset to capture diverse motions with detailed nuances, considering the physical aspects of the interaction. A large-scale 3D human motion dataset holds significant potential to catalyze novel research across various fields, including animation, gaming, virtual reality, and human-computer interaction. As of now, human motion datasets and synthesis methods have not resulted in any negative social impact.

### v2

@@ 5 Experiments
Figure 4: Nuanced interactions with the same object shape, but different resistance. a) The human struggles with the interaction as the resistance increases from 50 N to 200 N. At 500 N, the human fails to interact with the object. b) Plot depicts the horizontal shift of the center of mass of the human, estimated following [105, 14]. Results averaged over 1000 frames of the synthesized interaction.
@@ 5.1 Baselines
To our knowledge, the most related works to ours are NSM [98] and SAMP [33] as they both enable real-time generation human-object interaction with a user-controller. We also compare our method with InterDiff [120], a state-of-the-art method that leverages the diffusion model.
We train the baseline methods on our captured dataset using their publicly released codes. These methods neglect the physics attributes of the interaction. To measure the effectiveness of our intuitive physics encoding, we adapt NSM and SAMP by additionally conditioning on the resistive force and the binary hand contact. We refer to these modified baselines as NSM+ and SAMP+. Moreover, we evaluate two ablative variations of FORCE . One model is without the physics encoding $\mathcal{F}$, and the other one is without the contact prediction.
@@ 5.2 Quantitative Evaluation
Table 1: Quantitative comparisons for offline motion synthesis. Results evaluated on the testing sequences.
@@
NSM [98] 7.71 64.3 4.59
SAMP [33] 8.67 57.7 5.79
InterDiff [120] 8.62 75.3 6.82
NSM+ 7.08 69.6 4.71
SAMP+ 8.34 58.3 6.02
Ours no Phys. Enc. $\mathcal{F}$ 6.75 75.4 4.60
Ours no Contact Pred. - 77.2 4.48
Ours no R 7.10 76.4 4.76
Ours 6.02 84.0 4.46
Table 2: Quantitative comparisons for online full interaction synthesis. Note, diversity on the dataset is 0.906. The closeness to this figure indicates a better result.
@@
NSM [98] 0.707 75.1 5.94
SAMP [33] 0.719 66.2 7.72
NSM+ 0.785 79.6 6.01
SAMP+ 0.776 68.7 7.56
Ours no Phys. Enc. $\mathcal{F}$ 0.847 86.7 5.81
Ours no Contact Pred. 0.886 94.2 5.50
Ours 0.891 97.5 5.42
We quantitatively evaluate our model performance with two setups. a) offline motion synthesis. This is performed on the testing sequences given the ground truth control signals. We evaluate the accuracy (MPJPE) and motion quality (foot-slide and collision) of motion synthesis on testing sequences following our training/testing split explained in Section 4. The model synthesizes poses auto-regressively given the trajectories. The sequences are each distinctive, covering diverse pushing, pulling and carrying motions at different resistances. b) online full synthesis. We synthesize full approaching and interacting in a random setup. For fairness, 10 objects are randomly sampled. They are placed randomly relative to the human (within 6 meters) at a random orientation. The synthesis is performed 120 times for each method, repeating each action type (pushing, pulling, one-handed, and two-handed carrying) 30 times. We measure the diversity, the success rate and the execution time. Note, InterDiff [120] does not allow online synthesis, we compare against it for offline synthesis on the testing set.
@@
Foot-slide(cm) [40, 98, 36, 33, 136, 137]. The average traveled distance by the pivotal foot per step is calculated.
Diversity [32, 128, 138, 33]. We compute the Average Pairwise Distance (APD) on the root-relative joint positions.
[…]
Execution time [35, 80]. Time till a successful interaction.
From Table 1, it can be seen that  FORCE attains the lowest MPJPE of $6.02$ cm, surpassing both baseline models and ablative variations of FORCE . This emphasizes the crucial role of our intuitive physics encoding in facilitating the learning of multi-class motions. Additionally, our method minimizes object interpenetration, as reflected in the highest non-collision scores. It is notable that, compared with the ablative baseline without contact prediction, our method has a higher non-collision score. This indicates that with contact prediction, our model synthesizes more physically plausible motion. Furthermore, our model demonstrates improved motion quality, evident in its reduced foot-slide.
From Table 2, it can be seen under the online synthesis setup, FORCE has the highest success rate of interaction of $97.5\%$ and the shortest average execution time of $5.42$ seconds, illustrating the better controllability of our approach. Our model synthesizes motion of diversity (0.891) which is the closest match to the training data distribution (0.906). This underscores the significance of the novel physics encoding in enabling the synthesis of diverse, nuanced interactions.
@@ 5.3 Qualitative Evaluation
Figure 6: Without geometry representation $\mathbf{G}$ (left) vs. FORCE (right). The feature enables the reasoning about collision and avoid interpenetration.
@@ 5.4 Ablations
Resistance $\mathcal{R}$. As shown in table 1, without conditioning on the resistance, the motion quality degrades.
@@
Figure 7: a) FORCE predicts different hand contacts at different resistances for the same pushing motion. b) No contact prediction (left) vs. FORCE (right). FORCE exhibits less penetration.
@@
5.5 Generalization to shapes and locations.
As illustrated in Figure 8, FORCE generalizes to different shapes. Thanks to the object shape augmentation introduced in Section 4, our model is able to synthesize lifting a chair, an object category that is not seen during training of carrying sequences. It is also highlighted in Figure 8 that FORCE interacts with objects at different locations, for example, picking up from different heights.
Figure 8: (Top) FORCE generalizes to unseen shapes, demonstrated via pushing, carrying, and pulling. (Bottom) FORCE generalizes to different locations, performing one-handed and two-handed carry from the table and from the floor.
@@ 6 Conclusion
This work, FORCE , tackles the problem of synthesizing human object interactions with nuanced details, by modeling intuitive physics. It solves two main challenges. First, it proposes the first kinematic-based method that synthesizes human-object interaction conditioned on physical attributes such as object resistance and human force. By leveraging the novel physics encoding, our method generates a diverse spectrum of interaction motions under varying resistance. Its performance in diversity and realism surpasses previous methods. Our accompanying dataset features over 450 motion sequences. Together, they provide a benchmark for training and evaluating HOI methods.
@@ Acknowledgement
1 Failure Case
Figure 9: The interaction exhibits artifact when the object shape is too large.
The object augmentation that we employed for training enables FORCE to generalize to unseen object shapes at testing (see all the animations in the supplementary video.) However, when the object shape is too large, there may exist interpenetration artifact.
@@ 2 Dataset
Details on the dataset: Our dataset comprises 450 motion sequences involving human-object interactions with a diverse range of resistance forces. Table 3 provides a detailed breakdown of the dataset categorized by the level of resistance. In this context, resistance is measured solely by the mass of the removable weight used. The masses of the objects themselves are not factored into these measurements. They are measured and will be provided with the dataset. The dataset distributions based on the type of action (Table 4), and the type of hand contact (Table 5) are also presented.
@@
Table 5: Distribution of the dataset with different hand contact.
@@ 3 Architecture and Training Details
4 Architecture and Training Details